Bandar Industries Berhad of Malaysia manufactures sporting equipment. One of the company’s products, a football helmet for the North American market, requires a special plastic. During the quarter ending June 30, the company manufactured 3,400 helmets, using 2,176 kilograms of plastic. The plastic cost the company $16,538.

According to the standard cost card, each helmet should require 0.57 kilograms of plastic, at a cost of $8.00 per kilogram.

Required:

1. What is the standard quantity of kilograms of plastic (SQ) that is allowed to make 3,400 helmets?

2. What is the standard materials cost allowed (SQ × SP) to make 3,400 helmets?

3. What is the materials spending variance?

4. What is the materials price variance and the materials quantity variance?

(For requirements 3 and 4, indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values. Do not round intermediate calculations.)
Answer #1

1) Standard quality allowed = 3400*.57 = 1938 kg

2) Standard material cost = 1938*8 = $15504

3) Material spending variance = 15504-16538 = 1034 U

4) Material price variance = (8*2176-16538) = 870 F

Material quantity variance = (1938-2176)*8 = 1904 U